# E-Graphs as Circuits, and Optimal Extraction via Treewidth

### v1

@@ Abstract
We solve the optimal extraction problem for e-graphs by first showing a connection between e-graphs and cyclic monotone Boolean circuits, then solving the weighted satisfiability problem for such circuits. The solution is a parameterized algorithm based on treewidth. Additionally, we show how the circuit view of e-graphs allows us to apply simplification techniques that are not possible when operating directly on e-graphs. While the core parameterized algorithm may be adapted to work directly on e-graphs, the simplification results show why the circuit view is helpful.
@@ 1 Introduction
E-graphs are a type of directed graph with an equivalence relation on its nodes, and can be used to compactly represent exponentially many equivalent expressions. In recent years, this capability has found e-graphs many applications in formal methods, compilers, and automated reasoning communities.
@@
Common algorithmic techniques for NP-hard optimization problems include approximation algorithms and parameterized algorithms. It turns out that extraction is also hard to approximate to any constant factor [GLP24], so it is natural to turn to parameterized algorithms.
@@
    E-graphs that appear in practice often have low treewidth, a measure of how “close” to a tree a graph is. We quantify this in Section 4.2. Treewidth is commonly used in parameterized algorithms [Bod06].
@@
    E-graphs may be considered a certain class of monotone circuits, meaning a Boolean circuit with only AND and OR gates (no NOT gates). This allows us to draw on existing literature about treewidth algorithms for circuits to solve extraction.
@@
Our main contribution is the connection between circuits and e-graphs, and how that not only allows us to use existing circuit algorithms, but employ simplification techniques more effectively. Given that our main algorithm is extremely similar to the one given in [GLP24], we will omit giving our own practical implementation, evaluation against ILP methods, and proofs of correctness for the main algorithm, and direct the interested reader to their paper. While we do have a publicly available implementation, our focus is on the circuit connection and simplification.
@@ 2 Circuits and e-graphs
A weighted cyclic monotone circuit (henceforth “circuit”) is a directed graph with additional information $G=(V,E,V_{out},g,c)$. The set $V_{in}\subseteq V$ is the set of inputs, defined to be the set of vertices with in-degree 0. The set $V_{out}\subseteq V$ is the set of outputs. (These are the nodes whose values we are interested in, which might not have out-degree 0.) Finally, $g:V\setminus V_{in}\to\{\text{AND},\text{OR}\}$ is the gate type function and $c:V_{in}\to\mathbb{R}$ is a cost function. For any $U\subseteq V_{in}$, we denote $c(U)=\sum_{u\in U}c(u)$.
@@
A function $\alpha:V\to\{0,1\}$ is a (valid total) evaluation of $G$ if for all $u\in V$, denoting the inputs to $u$ as $\{v_{1},\dots,v_{k}\}$, we have $\alpha(u)=g(u)(\alpha(v_{1}),\dots,\alpha(v_{k})$. The evaluation satisfies $G$ if $\alpha(u)=1$ for all $u\in V_{out}$. It minimally satisfies $G$ if for every proper subset $A\subsetneq\{u\in V:\alpha(u)=1\}$, the function
@@
Let $\beta$ be an acyclic evaluation which agrees with $\alpha$ on $V_{in}$, we will show that $\alpha(u)=1$ iff $\beta(u)=1$. The forward and backward directions are identical, let us treat the forward direction.
Recall that $\alpha(u)=1$ iff $u\in G[\alpha]$. Because $G[\alpha]$ is acyclic, take the vertices in topological order. The base cases are the elements of $G[\alpha]$ with in-degree 0, note that these must have in-degree 0 in $G$ because gates cannot be 1 without at least one 1 input. Hence $\beta$ agrees with $\alpha$ here by hypothesis.
@@
Our main observation is that every e-graph can be represented as a circuit, in such a way that their semantics are equivalent. For an example, see Fig. 2.
Figure 2: An example converting e-graphs into circuits. Note that the arrows conventionally point to dependencies in e-graphs, but signals flow in the opposite direction in a circuit, so we flip the arrows. Furthermore, the extraction $A\mapsto\sqrt{\phantom{x}}$, $B\mapsto 2$ corresponds to the evaluation where everything on the left and the OR gate $A$ are all 1, and the rest are 0. The cyclic extraction $A\mapsto+$, $C\mapsto 0$ corresponds to the cyclic evaluation where everything on the right and the OR gate $A$ are all 1, and the rest are 0. The evaluation where everything is 1 has no corresponding extraction because $A$ can only choose one e-node in an e-graph, however such an evaluation is not minimal.
@@
Then there exists an acyclicity-preserving bijection between minimal satisfying extractions of $\mathcal{G}$ and minimal satisfying evaluations of $G$.
The proof is a tedious checking of these definitions that does not require any external results, the details are contained in Appendix A. With this observation, in order to solve the extraction problem for e-graphs, it suffices to solve the weighted satisfiability problem for (potentially cyclic) circuits.
@@ 3.1 Preliminaries
Our main inspiration is Proposition 4.7 of [KTX17], which solved the weighted minimum satisfiability problem for acyclic monotone circuits by parameterizing on treewidth. Based on the reduction illustrated in the previous section, the only additional algorithmic contribution that we need to make is to describe why the cyclic nature of the graph is not a problem, and how we enforce the extraction result to be acyclic. For completeness, we will recap the full algorithm (with slightly different notation from the original paper).
The main technique is called treewidth, or tree decomposition. This is a classical technique, see Chapter 7 of [CFK+15] for more in-depth information.
@@
    For all $v\in V$, the subgraph of $\mathcal{T}$ induced by the bags which contain $v$ form a tree.
The width of a tree decomposotion is the size of the largest bag, minus one. The treewidth of a graph is the smallest width of any tree decomposition. If $\mathcal{T}$ is rooted, we write $T_{X}$ for the union of all bags underneath $X\in\mathcal{X}$ (including $X$). A nice tree decomposition is one in which every bag $X$ is one of 4 kinds:
@@
The core feature of tree decomposition is that as you walk up the tree, the current bag cuts the original graph into two disconnected pieces, the set of vertices that you have already seen and forgotten, and the set of vertices that you have not seen yet. This property allows us to do dynamic programming and focus only on the current bag.
One last definition that we will use in the algorithm is that of a partial evaluation, since we would like to incrementally build up our evaluations at each bag. We will use the following definition with $U=T_{X}$.
@@
A (locally valid) partial evaluation of a circuit is a function $\alpha:U\to\{0,1\}$ where $U\subseteq V$, and for all $u\in U$, denoting the inputs to $u$ as $\{v_{1},\dots,v_{k},w_{1},\dots,w_{\ell}\}$ where each $v_{i}\in U$ and $w_{i}\not\in U$, there exist $b_{1},\dots,b_{\ell}\in\{0,1\}$ such that $\alpha(u)$ is $g(u)$ (AND or OR) applied to $(\alpha(v_{1}),\dots,\alpha(v_{k}),b_{1},\dots,b_{\ell})$.
@@ 3.2 The algorithm
    We do dynamic programming. At every bag, for every possible summary of a partial evaluation, the program will remember the minimum cost partial evaluation producing that summary.
    The summary of a partial evaluation $\alpha:T_{X}\to\{0,1\}$ at bag $X$ has three parts. First, there is the restriction $\alpha|_{X}$. Second, there is the map $\text{known}_{\alpha}|_{X}$, where $\text{known}_{\alpha}(u)=1$ if and only if $\alpha(u)=1$ and that this fact can be deduced from the value of $\alpha$ on the inputs to $u$. (For example, if $u$ is an AND gate and one of its inputs is not in $\operatorname{dom}(\alpha)$, then $\text{known}_{\alpha}(u)=0$.) Lastly, there is $G[\alpha]^{+}[X]$, the subgraph of the transitive closure of $G[\alpha]$ induced by $X$. In other words, which elements in $X$ have paths between them in $G[\alpha]$? This is what allows us to correctly handle acyclicity, remembering which vertices in the bag are connected, even when the connections themselves have already been forgotten from the bag, and is the only change from [KTX17].
@@ 4.1 Rules
One natural way to improve the speed of our algorithm is to simplify the instances directly. We found in our testing that this helps dramatically.
Finding the most compact representation of a Boolean circuit is often known as circuit minimization. Circuit minimization is well-known to be NP-hard, but there is an abundance of existing software to minimize at least to a best effort, such as Espresso [BSVMH84]. However, because our circuits are cyclic, and we also want to use properties of extraction to make simplifications beyond logical ones, we implemented several of our own heuristic rules.
For all of the rules below, let $G=(V,E,V_{out},g,c)$ be a monotone circuit. These rules are by no means an exhaustive list of all possible simplifications, they are only an exploratory list of the kinds of rules that may be beneficial.
@@
    One may check for both AND and OR that after deleting the edge, the dependency still is maintained through the path.
@@
    Suppose $v_{n}$ is true. Then by induction, every $v_{i}$ is true, as well as $u$. Then this evaluation has a true cycle. So the optimal acyclic evaluation must have $v_{n}$ false, so we can delete it.
@@
Some of these rules are purely based on the circuit structure (rules 2, 3, 4, 5). These were rules that we happened to try by chance, but in future work, one might consider taking an existing general purpose circuit minimization tool and adapting all parts compatible with cyclic circuits. The great benefit of representing e-graphs as circuits is that it unlocks a plethora of existing work to learn from.
Other rules are more specific to extraction (rules 1, 6, 7). We note that some rules can be generalized to all extraction algorithms, for instance, rule 1. However, many combinations of these rules lead to situations that e-graphs are too inflexible for. Recall that the main algorithm works as long as the circuit is monotone, and that direct translations of e-graphs always result in an alternating AND/OR pattern with every AND gate having outdegree 1. Therefore, the circuit view also helps us leverage deeper and more complicated simplifications.
@@ 4.2 Simplification evaluation
Unfortunately, we found that the vast majority of e-graphs in this test set were too large for exact tree decomposition algorithms to work in a reasonable amount of time. However, approximate tree decomposition algorithms suffice. We used the “arboretum” Rust library, which implements various heuristics but primarily relies on the classical minimum degree heuristic to compute an upper bound on treewidth. Note that this is not an issue for the main algorithm’s correctness, which only relies on having a valid tree decomposition of any width—a larger width only affects the running time.
Although the upper bound is not an exact measure, it can still give a rough idea of how our simplification schemes affect treewidth. It is also a realistic measure, since the important quantity is the width of the tree decomposition available to our algorithm, not the true treewidth. In our implementation, we apply each of the rules in a loop until we reach a fixed point. The effect of simplification on treewidth and $|V|$ is shown in Fig. 5, and these effects are quantified in Table 2.
@@
It is also notable that every collection of graphs in our dataset exhibits a linear relationship between $|V|$ and treewidth. This may be a consequence of the fact that each source generally produces graphs of similar average degree. The behavior is reminiscent of that for random graphs: in a random graph where each edge has probability $c/n$ of appearing for $c>1$, the treewidth of the graph is $\Omega(n)$ [LLO11] and upper bounded by $tn$ for some $t<1$ [WLCX11], the exact behavior that we see here. This suggests that sparsity is the only contributing factor to low treewidth in this data, not any deeper features of the e-graph generation process.
@@
Lastly, we note [GLP24] claims that their implementation is faster than ILP solvers for most e-graphs with treewidth under 10. Few e-graphs in our test set had treewidth under 10 before simplification, but especially for the egraphs in the collections “eggcc-bril” and “fuzz”, a significant fraction of them have treewidth under 10 after simplification, bringing them within the range where treewidth methods already surpass other existing methods.
@@ 5 Future directions
A second open direction is to investigate more closely which methods of e-graph creation lead to smaller treewidths, and then design e-graph creation methods (or saturation, as it is often called) that minimize treewidth from the start. Since treewidth is often calculated with heuristic algorithms, saturation methods could even be designed with a particular treewidth heuristic in mind to further improve efficiency.
@@
  • To show that $\alpha$ is minimally satisfying, it is clearly satisfying, so it remains to show minimality. Suppose for contradiction that there exists a proper subset $A\subsetneq\{i\in V:\alpha(u)=1\}$ such that $\alpha|_{A}$ is a satisfying evaluation.
@@
    First, note that $\mathcal{A}$ is a subset of $\operatorname{dom}(\varphi)$ because $\alpha|_{A}(\lor_{C})=1$ implies $C\in\operatorname{dom}(\varphi)$ by ($\dagger$). Next, we show that it is a proper subset. Because $A$ is a proper subset of $\{i\in V:\alpha(i)=1\}$, we get $i\in V\setminus A$ with $\alpha(i)=1$.
@@
      If $i=\lor_{C}$, then $\alpha(\lor_{C})=1$ means we have found $C\in\operatorname{dom}(\varphi)\setminus\mathcal{A}$ as desired.
@@
      If $i=\land_{u}$, then $\alpha(\land_{u})\neq\alpha|_{A}(\land_{u})$, we note that by our construction of $\alpha$, if $C$ is the class containing $u$, then $\land_{u}$ is the only true input to $\lor_{C}$ in $\alpha$. Now with $\alpha|_{A}(\land_{u})=0$, because $\alpha|_{A}$ is a valid evaluation, we must have case (1).
@@
      If $i=x_{u}$, then $\alpha(x_{u})\neq\alpha|_{A}(x_{u})$, again by our construction of $\alpha$ and the fact that $\alpha|_{A}$ is a valid evaluation, we get $\alpha|_{A}(\land_{u})=0$, and we are in case (2).
    Now to show that $\varphi|_{\mathcal{A}}$ is a satisfying extraction. First, to show that it is a valid extraction, whenever $C\in\mathcal{A}$, because only one input to $\lor_{C}$ was true in $\alpha$, namely $\land_{\varphi(C)}$, it must be that $\alpha|_{A}(\land_{\varphi(C)})=1$. If $\varphi(C)$ depends on $C_{1},\dots,C_{k}$, it must be that $\alpha|_{A}(\lor_{C_{i}})=1$ for each $i$, so $C_{i}\in\mathcal{A}$ as required for an extraction. The extraction is satisfying because $\alpha|_{A}$ being satisfying implies $\alpha|_{A}(\lor_{C})=1$ for all $C\in\mathcal{C}_{out}$, and hence $C\in\mathcal{A}$ as required.
@@
  • To show that $\varphi$ is an extraction, simply note that whenever $C\in\operatorname{dom}(\varphi)$ is defined, there exists $u\in C$ with $\alpha(\land_{u})=1$, so all dependencies $C_{i}$ must have $\alpha(\lor_{C_{i}})=1$. This means that for each $C_{i}$, at least one $u_{i}\in C_{i}$ must have $\alpha(\land_{u_{i}})=1$, so $C_{i}\in\operatorname{dom}(\varphi)$ and we are done.
  • To show that $\varphi$ is satisfying, simply note that because $\alpha$ is satisfying, $\alpha(\land_{out})=1$, which is only possible if $\alpha(\lor_{C})=1$ for all $C\in\mathcal{C}_{out}$. Each $\lor_{C}$ can only be 1 if at least one of its children $\land_{u}$ is evaluated to 1 where $u\in C$, so $C\in\operatorname{dom}(\varphi)$ and we are done.
@@
Lastly, to show that inverse map is truly an inverse, we need to show that transforming $\varphi$ to $\alpha$ to $\varphi$ is the identity, which is obvious, and that transforming $\alpha$ to $\varphi$ to $\alpha$ is identity, for which it suffices to note that $\alpha$ is entirely determined by its values on $\land_{u}$: vertices $\lor_{C}$ only have vertices of type $\land_{u}$ as inputs, so they are determined, likewise $\land_{out}$ which only has $\lor_{C}$ as inputs, and we must have $\alpha(x_{u})=\alpha(\land_{u})$, because $\alpha(x_{u})=0$ with $\alpha(\land_{u})=1$ is not valid and $\alpha(x_{u})=1$ with $\alpha(\land_{u})=0$ is not minimal (take $A=V\setminus\{x_{u}\}$).
@@
  • Suppose $\varphi$ is acyclic. To show that $G[\alpha]$ is acyclic, it suffices to show that every directed cycle in $G$ has at least one vertex on which $\alpha$ is 0. The only possible directed cycles of $G$ occur as alternations of edges of type $(\land_{u},\lor_{C})$ and $(\lor_{C},\land_{u})$. So let $\land_{u_{1}},\lor_{C_{1}},\dots,\land_{u_{k}},\lor_{C_{k}},\land_{u_{k+1}}=\land_{u_{1}}$ be a cycle in $G$, where $u_{i}\in C_{i}$ and $(C_{i},u_{i+1})\in\mathcal{E}$ for all $i$, and it would suffice to find $u_{i}$ such that $\alpha(\land_{u_{i}})=0$. Because $\varphi$ is acyclic, there must be some $u_{i}$ for which $\varphi(C_{i})\neq u_{i}$ (otherwise $C_{1},\dots,C_{k},C_{1}$ is a cycle). Therefore $\alpha(\land_{u_{i}})=0$ as desired.

### v2

@@ Abstract
We demonstrate a new connection between e-graphs and Boolean circuits. This allows us to adapt existing literature on circuits to easily arrive at an algorithm for optimal e-graph extraction, parameterized by treewidth, which runs in $2^{O(w^{2})}\text{poly}(w,n)$ time, where $w$ is the treewidth of the e-graph. Additionally, we show how the circuit view of e-graphs allows us to apply powerful simplification techniques, and we analyze a dataset of e-graphs to show that these techniques can reduce e-graph size and treewidth by 40-80% in many cases. While the core parameterized algorithm may be adapted to work directly on e-graphs, the primary value of the circuit view is in allowing the transfer of ideas from the well-established field of circuits to e-graphs.
@@ 1 Introduction
E-graphs are a type of directed graph with an equivalence relation on its nodes that can be used to compactly represent exponentially many equivalent expressions. In recent years, this capability has found e-graphs to have many applications in formal methods, compilers, and automated reasoning communities ([JNR02], [TSTL09], [STL11], [WNW+21]).
@@
Common algorithmic techniques for NP-hard optimization problems include approximation algorithms and parameterized algorithms. It turns out that extraction is also hard to approximate to any constant factor [GLP24], so it is natural to turn to parameterized algorithms: algorithms that are efficient after a particular parameter is chosen to be fixed. One commonly used parameter is treewidth, a measure of how “close” to a tree the graph is (a survey is available in [Bod06]).
@@
    E-graphs that appear in practice often have low treewidth, so it is a good choice for parameterization. We quantify this in Section 4.2.
@@
    E-graphs may be considered a certain class of monotone circuits, meaning a Boolean circuit with only AND and OR gates (no NOT gates). This allows us to draw on existing literature about treewidth-based algorithms for circuits in order to solve extraction.
@@
Our main contribution is the connection between circuits and e-graphs, which not only allows us to use existing circuit algorithms but also employ simplification techniques more effectively. Given that our main algorithm is extremely similar to the one given in [GLP24], we will omit to give our own practical implementation, evaluation against existing extraction methods, and proofs of correctness for the main algorithm, and direct the interested reader to their paper. While we do have a publicly available implementation [^2] https://github.com/glenn-sun/egg-extraction-gym/tree/glenn-treewidth/src/extract/treewidth, our focus is on the circuit connection and simplification.
@@ 2 Circuits and e-graphs
A weighted cyclic monotone circuit (henceforth “circuit”) is a directed graph with additional information $G=(V,E,V_{out},g,c)$. The set $V_{in}\subseteq V$ is the set of inputs, defined to be the set of vertices with in-degree 0. The set $V_{out}\subseteq V$ is the set of outputs. (These are the nodes whose values we are interested in, which may or may not have out-degree 0.) Finally, $g:V\setminus V_{in}\to\{\text{AND},\text{OR}\}$ is the gate type function and $c:V_{in}\to\mathbb{R}$ is a cost function. For any $U\subseteq V_{in}$, we denote $c(U)=\sum_{u\in U}c(u)$.
@@
A function $\alpha:V\to\{0,1\}$ is a (valid total) evaluation of $G$ if for all $u\in V$, denoting the inputs to $u$ as $\{v_{1},\dots,v_{k}\}$, we have $\alpha(u)=g(u)(\alpha(v_{1}),\dots,\alpha(v_{k}))$. The evaluation satisfies $G$ if $\alpha(u)=1$ for all $u\in V_{out}$. It minimally satisfies $G$ if for every proper subset $A\subsetneq\{u\in V:\alpha(u)=1\}$, the function
@@
Let $\beta$ be an acyclic evaluation which agrees with $\alpha$ on $V_{in}$, we will show that $\alpha(u)=1$ iff $\beta(u)=1$. The forward and backward directions are identical. Let us treat the forward direction.
Recall that $\alpha(u)=1$ iff $u\in G[\alpha]$. Because $G[\alpha]$ is acyclic, take the vertices in topological order. The base cases are the elements of $G[\alpha]$ with in-degree 0; note that these must have in-degree 0 in $G$ because gates cannot be 1 without at least one 1 input. Hence $\beta$ agrees with $\alpha$ here by hypothesis.
@@
Our main observation is that every e-graph can be represented as a circuit in such a way that its semantics are equivalent. For an example, see Fig. 2.
Figure 2: An example of converting e-graphs into circuits. Note that the arrows conventionally point to dependencies in e-graphs, but signals flow in the opposite direction in a circuit, so we flip the arrows. Furthermore, the extraction $A\mapsto\sqrt{\phantom{x}}$, $B\mapsto 2$ corresponds to the evaluation where everything on the left and the OR gate $A$ are all 1, and the rest are 0. The cyclic extraction $A\mapsto+$, $C\mapsto 0$ corresponds to the cyclic evaluation where everything on the right and the OR gate $A$ are all 1, and the rest are 0. The evaluation where everything is 1 has no corresponding extraction because $A$ can only choose one e-node in an e-graph; however, such an evaluation is not minimal.
@@
Then, there exists an acyclicity-preserving bijection between minimal satisfying extractions of $\mathcal{G}$ and minimal satisfying evaluations of $G$.
The proof is a tedious checking of these definitions that does not require any external results. The details are contained in Appendix A. With this observation, in order to solve the extraction problem for e-graphs, it suffices to solve the weighted satisfiability problem for (potentially cyclic) circuits.
@@ 3.1 Preliminaries
Our main inspiration is Proposition 4.7 of [KTX17], which solved the weighted minimum satisfiability problem for acyclic monotone circuits by parameterizing on treewidth. Based on the reduction illustrated in the previous section, the only additional algorithmic contribution that we need to make is to describe why the cyclic nature of the graph is not a problem and how we enforce the extraction result to be acyclic. For completeness, we will recap the full algorithm (with slightly different notations from the original paper).
The main technique is called treewidth, or tree decomposition. This is a classical technique; see Chapter 7 of [CFK+15] for more information.
@@
    For all $v\in V$, the subgraph of $\mathcal{T}$ induced by the bags that contain $v$ forms a tree.
The width of a tree decomposition is the size of the largest bag minus one. The treewidth of a graph is the smallest width of any tree decomposition. If $\mathcal{T}$ is rooted, we write $T_{X}$ for the union of all bags underneath $X\in\mathcal{X}$ (including $X$). A nice tree decomposition is one in which every bag $X$ is one of 4 kinds:
@@
The core feature of tree decomposition is that as you walk up the tree, the current bag cuts the original graph into two disconnected pieces: the set of vertices that you have already seen (including those that are forgotten) and the set of vertices that you have not seen yet. This property allows us to do dynamic programming and focus only on the current bag.
One last definition that we will use in the algorithm is that of a partial evaluation since we would like to build up our evaluations at each bag incrementally. We will use the following definition with $U=T_{X}$.
@@
A (locally valid) partial evaluation of a circuit is a function $\alpha:U\to\{0,1\}$ where $U\subseteq V$, and for all $u\in U$, denoting the inputs to $u$ as $\{v_{1},\dots,v_{k},w_{1},\dots,w_{\ell}\}$ where each $v_{i}\in U$ and each $w_{j}\not\in U$, there exist $b_{1},\dots,b_{\ell}\in\{0,1\}$ such that $\alpha(u)$ is $g(u)$ (AND or OR) applied to $(\alpha(v_{1}),\dots,\alpha(v_{k}),b_{1},\dots,b_{\ell})$.
@@ 3.2 The algorithm
    We do dynamic programming. At every bag, for every possible summary of a partial evaluation, the program will remember the minimum cost partial evaluation producing that summary. The summary of a partial evaluation $\alpha:T_{X}\to\{0,1\}$ at bag $X$ has three parts:
    • First, there is the restriction $\alpha|_{X}$.
    • Second, there is the map $\text{known}_{\alpha}|_{X}$, where $\text{known}_{\alpha}(u)=1$ if and only if $\alpha(u)=1$ and that this fact can be deduced from the value of $\alpha$ on the inputs to $u$. For example, if $u$ is an AND gate and one of its inputs is not in $\operatorname{dom}(\alpha)$, then $\text{known}_{\alpha}(u)=0$.
    • Lastly, there is $G[\alpha]^{+}[X]$, the subgraph of the transitive closure of $G[\alpha]$ induced by $X$. In other words, which elements in $X$ have paths between them in $G[\alpha]$? This is what allows us to correctly handle acyclicity, remembering which vertices in the bag are connected, even when the connections themselves have already been forgotten from the bag, and is the only change from [KTX17].
    We break into cases depending on the type of bag.
@@ 4.1 Rules
One natural way to improve the speed of our algorithm is to simplify the instances directly. We found in our testing that this helps dramatically. The idea of preprocessing e-graphs to simplify them before extraction is not new (see for example, the repository at [Han24]), but the circuit view of e-graphs makes this process easier, more transparent, and more powerful.
Finding the most compact representation of a Boolean circuit is often known as circuit minimization. Circuit minimization is well-known to be NP-hard, but there is an abundance of existing software to quickly attempt a best effort, such as SIS [SSL+92] among others.
Note that in order to use off-the-shelf software, it must support cycles. Though our particular semantics are unique, all simplifications valid for sequential circuits are valid for us, a type of circuits fundamental in hardware. Sequential circuits are cyclic circuits where the “undefined” behavior is explicitly specified by propagation delay and keeping track of the circuit’s state. A minimizer for sequential circuits will additionally ensure that these stateful behaviors are preserved, which we can just ignore—it must also preserve behaviors independent of state, as acyclic extractions are. Thus, we may want to do some additional simplification afterwards specific to our own extraction semantics, but such software may be used as a first step.
To show some basic ideas, as well as to highlight properties of extraction to make simplifications beyond what is possible from off-the-shelf software, we implemented several of our own heuristic rules. For all of the rules below, let $G=(V,E,V_{out},g,c)$ be a monotone circuit. These rules are by no means an exhaustive list of all possible simplifications, they are only an exploratory list of the kinds of rules that may be beneficial.
@@
    One may check for both AND and OR that after deleting the edge, the dependency is still maintained through the path.
@@
    Suppose $v_{n}$ is true. Then by induction, every $v_{i}$ is true, as well as $u$. Then, this evaluation has a true cycle. So, the optimal acyclic evaluation must have $v_{n}$ false, and we can delete it.
@@
Some of these rules are purely based on the circuit structure (rules 2, 3, 4, 5). These were rules that we arbitrarily decided to implement, but in future work, one might consider replacing with an existing general purpose circuit minimization tool. This is one potential benefit of circuits that we have yet to explore. Other rules are more specific to extraction (rules 1, 6, 7). We note that some rules can be generalized to all extraction algorithms, such as rule 1.
Although there are often equivalent rules that operate on e-graphs directly without translating to circuits, circuits can generally be smaller (not to mention unlocking the ability to harness existing software). Recall that a direct translations of an e-graph will always result in an alternating AND/OR pattern, with every AND gate having outdegree 1 and one variable as an in-neighbor. Rules like rule 2 and rule 7 break this structure, and they are valid because the algorithm works for all monotone circuits.
@@ 4.2 Simplification evaluation
Note that the main algorithm’s correctness only relies on having a valid tree decomposition of any width—a larger width only affects the running time. Thus, although we unfortunately found that the vast majority of e-graphs in this test set were too large for computing exact tree decomposition, approximate tree decomposition algorithms suffice. We used the “arboretum” Rust library, which implements various heuristics but primarily relies on the classical minimum degree heuristic to compute an upper bound on treewidth.
In order to compare the effects of our simplification scheme, such an upper bound is also a more realistic measure to compare than the true treewidth, since the important quantity is the width of the tree decomposition available to our algorithm. In our implementation, we apply each of the rules in a loop until we reach a fixed point. The effect of simplification on treewidth and $|V|$ is shown in Fig. 5, and these effects are quantified in Table 2.
@@
We note that although the treewidth of most e-graphs in our dataset is small, sparsity is likely to be the only contributing factor to low treewidth in these data, not any deeper features of the e-graph generation process. This is because the linear relationship between $|V|$ and treewidth is the exact relationship predicted by random graphs of constant average degree: In a random graph where each edge has probability $c/n$ of appearing for $c>1$, the treewidth of the graph is $\Omega(n)$ [LLO11] and upper bounded by $tn$ for some $t<1$ [WLCX11].
@@
Lastly, we note that these simplification methods bring many more e-graphs into the range in which treewidth-based methods are faster than existing methods. Though we did not develop an efficient implementation ourselves, [GLP24] claims that their implementation of treewidth-based extraction is faster than ILP solvers for most e-graphs with treewidth under 10, at least when not required to output acyclic extractions. Few e-graphs in our test set had treewidth under 10 before simplification, but especially for the egraphs in the collections “eggcc-bril” and “fuzz”, a significant fraction of them have treewidth under 10 after simplification.
@@ 5 Future directions
A second open direction is to investigate more closely if particular methods of e-graph creation lead to smaller treewidth, and then design e-graph creation methods (or saturation, as it is often called) that minimize treewidth from the start. Since treewidth is often calculated with heuristic algorithms, it would be interesting to determine if certain saturation methods could even be designed with a particular treewidth heuristic in mind to further improve efficiency.
@@
  • To show that $\alpha$ is minimally satisfying, it is certainly satisfying from the definitions of $\alpha$ and $V_{out}$, so it remains to show minimality. Suppose for contradiction that there exists a proper subset $A\subsetneq\{i\in V:\alpha(i)=1\}$ such that $\alpha|_{A}$ is a satisfying evaluation.
@@
    First, note that $\mathcal{A}$ is a subset of $\operatorname{dom}(\varphi)$ because $\alpha|_{A}(\lor_{C})=1$ implies $C\in\operatorname{dom}(\varphi)$ by ($\dagger$). Next, we show that it is a proper subset. Because $A$ is a proper subset of $\{i\in V:\alpha(i)=1\}$, there is a true vertex (by $\alpha$) outside of $A$.
@@
      If it is $\lor_{C}$, then $\alpha|_{A}(\lor_{C})=1$ and $C$ is a satisfied class outside of $\mathcal{A}$.
@@
      If it is $\land_{u}$, by our construction ($\dagger$), no siblings of $\land_{u}$ are true. So if $\land_{u}\not\in A$ and $C$ denotes the class containing $u$, by validity of $\alpha|_{A}$, we have $\alpha|_{A}(\lor_{C})=0$, and $C$ is a satisfied class outside of $\mathcal{A}$.
@@
      If it is $x_{u}$, by validity of $\alpha|_{A}$, the vertex $\land_{u}$ must be false in $\alpha|_{A}$, and $\land_{u}$ must be true in $\alpha$ by ($\dagger$). Therefore, we are in case (2) and can repeat the argument to find the desired $C$.
    Now to show that $\varphi|_{\mathcal{A}}$ is a satisfying extraction. First, to show that it is a valid extraction, if $\varphi(C)$ depends on $C_{1},\dots,C_{k}$, by validity of $\alpha|_{A}$ it must be that $\alpha|_{A}(\lor_{C_{i}})=1$ for each $i$, so $C_{i}\in\mathcal{A}$ as required for an extraction. The extraction is satisfying because $\alpha|_{A}$ being satisfying implies $\alpha|_{A}(\lor_{C})=1$ for all $C\in\mathcal{C}_{out}$, and hence $C\in\mathcal{A}$ as required.
@@
  • To show that $\varphi$ is an extraction, simply note that whenever $\varphi(C)$ is defined, by the above construction $\alpha(\land_{\varphi(C)})=1$, so by validity of $\alpha$ and ($\star$), all dependencies $C_{i}$ must have $\alpha(\lor_{C_{i}})=1$. Again by validity, this means that for each $C_{i}$, at least one $u_{i}\in C_{i}$ must have $\alpha(\land_{u_{i}})=1$, so $C_{i}\in\operatorname{dom}(\varphi)$ and we are done.
  • To show that $\varphi$ is satisfying, simply note that because $\alpha$ is satisfying, $\alpha(\lor_{C})=1$ for all $C\in\mathcal{C}_{out}$. Each $\lor_{C}$ can only be 1 if at least one of its children $\land_{u}$ is evaluated to 1 where $u\in C$, so $C\in\operatorname{dom}(\varphi)$ and we are done.
@@
Lastly, to show that inverse map is truly an inverse, we need to show that transforming $\varphi$ to $\alpha$ to $\varphi$ is the identity, which is obvious, and that transforming $\alpha$ to $\varphi$ to $\alpha$ is identity, for which it suffices to note that $\alpha$ is entirely determined by its values on $\land_{u}$: vertices $\lor_{C}$ only have vertices of type $\land_{u}$ as inputs, so they are determined, and we must have $\alpha(x_{u})=\alpha(\land_{u})$, because $\alpha(x_{u})=0$ with $\alpha(\land_{u})=1$ is not valid and $\alpha(x_{u})=1$ with $\alpha(\land_{u})=0$ is not minimal (take $A=V\setminus\{x_{u}\}$).
@@
  • Suppose $\varphi$ is acyclic. To show that $G[\alpha]$ is acyclic, it suffices to show that every directed cycle in $G$ has at least one vertex on which $\alpha$ is 0. The only possible directed cycles of $G$ occur as alternations of edges of type $(\land_{u},\lor_{C})$ and $(\lor_{C},\land_{u})$. So let $\land_{u_{1}},\lor_{C_{1}},\dots,\land_{u_{k}},\lor_{C_{k}},\land_{u_{k+1}}=\land_{u_{1}}$ be a cycle in $G$, where $u_{i}\in C_{i}$ and $(u_{i+1},C_{i})\in\mathcal{E}$ for all $i$, and it would suffice to find $u_{i}$ such that $\alpha(\land_{u_{i}})=0$. Because $\varphi$ is acyclic, there must be some $u_{i}$ for which $\varphi(C_{i})\neq u_{i}$ (otherwise $C_{1},\dots,C_{k},C_{1}$ is a cycle). Therefore by ($\dagger$), $\alpha(\land_{u_{i}})=0$ as desired.